# LSSInst: Improving Geometric Modeling in LSS-Based BEV Perception with Instance Representation

### v1

@@ 3 Methodology
Figure 2: Overview of LSSInst. The multi-view images with previous $T$ frames are fed into the backbone network for the image features. BEV branch looks around the image feature to generate the BEV feature by view transformation and temporal encoding. Instance adaptor aggregates the sparse object-aware feature from the BEV feature and prepares the multiplicate 3D query combination. Instance branch looks back at the image feature and perfects the sparse feature by spatiotemporal sampling and fusion. Lastly, the model makes the final prediction based on the updated output.
Utilizing instance-level representations on the basis of the scene-level BEV to excavate more detailed features and geometric information is of practical significance for generalized 3D perception. In this work, we propose LSSInst, which looks back for the more geometry-aware and fine-grained target feature extraction to bridge the adaptation between scene-level and instance-level 3D representations. The overview of our framework is shown in Fig. 2, and we organize the remaining part as follows. Firstly, Sec. 3.1 briefly introduces the BEV branch. Next, Sec. 3.2 introduces the instance adaptor module, and the instance branch is given in Sec. 3.3.
@@ 3.1 BEV Branch: Looking around for scene-level representation
The multi-view sequential images with the previous $T$ frames are first input into the 2D image backbone network for feature extraction. Then the BEV branch receives the extracted image feature $F_{img}\in\mathbb{R}^{N_{v}\times N_{s}\times C\times H\times W}$ and functions as an around looker, translating $F_{img}$ from 2D camera views to BEV for preliminary scene-level representation $F_{\beta}\in\mathbb{R}^{C\times H_{\beta}\times W_{\beta}}$, where $N_{v}$, $N_{s}$ denote the camera-view and scale number respectively. This branch can be briefly divided into temporally-shared view transformation for BEV generation and BEV sequence fusion. The 2D-to-BEV view transformation (VT) is naturally based on LSS paradigm, which can be mainly concluded as the depth refinement module (DRM), feature extraction net, and voxel pooling. For the best version of the framework, we adopted the Gaussian-spaced top-k stereo [34] for a better depth distribution map before voxel pooling. After the shared VT, a sequence of BEV representations will be aligned into current time $t$ and fed to the BEV temporal encoder to form the final current BEV. Here the encoder is designed as a very lightweight residual network for dimension reduction only.
3.2 Instance Adaptor: Scene-to-instance adaptation
For the sake of preserving a coherent and solid semantic consistency between BEV and instance representations, we propose the instance adaptor module to eliminate the gaps in the position description and space discrepancy. Since the BEV feature $F_{\beta}$ is a scene-level representation surrounding the ego car, there is redundancy and inflexibility in modeling instance-level features. To that end, the proposed adaptor module first performs a reprojection of the proposal box coordinates $P_{o}\in\mathbb{R}^{N_{\beta}\times 3}$ obtained through the BEV proposal head, returning to the BEV-recognized position $P_{\beta}\in\mathbb{R}^{N_{\beta}\times 2}$ to resample the object-related features. Here $N_{\beta}$ denotes the number of BEV proposals. Given the BEV point-cloud range $R_{\beta}$ with the corresponding voxel size $S_{v}$ and up-sampling factor $\sigma$, we can formulate the 2D reprojected coordinate as follow:
@@ 3.2 Instance Adaptor: Scene-to-instance adaptation
here $z_{\beta}$ denotes the z-axis homogeneous term and is actually a constant of 1. Moreover, due to the overfitting bias in BEV, the focused area may deviate from the actual object position. Inspired by deformable attention [56], the adaptor module incorporates the learnable offsets based on the original focused feature $F_{\beta}^{{}^{\prime}}\in\mathbb{R}^{N_{\beta}\times C}$ for misalignment compensation by exploring more semantically-aware regions. Suppose $i\in\{1,2,...,N_{\beta}\}$ denotes an arbitrary element index of $F_{\beta}^{{}^{\prime}}$ and its aggregated instance-wise feature $F_{\alpha}^{i}$ can be formulated by
@@ 3.3 Instance Branch: Looking back for instance-level representation
Given the sequential image features $\{F_{img}^{t}\}_{t=0}^{T_{\chi}}$ $(T_{\chi}\leq T)$ from the image backbone network and the sparse instance features $F_{\chi}$ with corresponding 3D boxes $P_{box}\in\mathbb{R}^{N\times C_{box}}$ from the instance adaptor, the instance branch will spatially and temporally look back the image feature based on the referenced box coordinates and iteratively extract the abundant but more fine-grained representations to update pre-features. This branch can be roughly regarded as a multi-layer Transformer-decoder-like [42] module for 3D detection, which is briefly divided into two parts: box-level offset and embedding, as well as spatiotemporal sampling and fusion.
@@
Different from the previous DETR-like 3D approaches such as DETR3D [46], Polarformer [20], VEDet [3] that iteratively refine just with 3D coordinate offset regression, the instance branch adopts the box-level offset regression based on $P_{box}$. With this convenience, we can encode all the geometric-aware information of the entire box to substitute the transitional positional encoding, thereby expanding and enriching the space of feature expression rather than the superficial positional level. With it combined with the sparse instance features, there will be more geometric priors and implicit compensation in subsequent attention interactions. Precisely, we first categorize $P_{box}$ based on the element semantic of box dimension into four divisions, which are position $P_{pos}\in\mathbb{R}^{N\times 3}$ (i.e., $x,y,z$), scale $P_{sca}\in\mathbb{R}^{N\times 3}$ (i.e., $w,l,h$), velocity $P_{vel}\in\mathbb{R}^{N\times 2}$ (i.e., $v_{x},v_{y}$), and orientation $P_{ori}\in\mathbb{R}^{N\times 2}$ (i.e., $sin(\theta_{yaw}),cos(\theta_{yaw})$) respectively. Then we introduce five separated linear projections $\{\textbf{E}_{l3}^{j}\}^{2}_{j=1}\in\mathbb{R}^{3\times C}$, $\{\textbf{E}_{l2}^{j}\}^{2}_{j=1}\in\mathbb{R}^{2\times C}$ and $\textbf{E}_{g}\in\mathbb{R}^{C\times C}$ for comprehensive encoding, of which the former four embed every category locally and the last one embeds them globally. The final box embedding $G_{\chi}\in\mathbb{R}^{N\times C}$ can be formulated by
@@
The sparse feature $F_{\chi}$ with the box embedding $G_{\chi}$ will be updated by the spatial and temporal sampling after being fed into the Multi-Head Self-Attention block [42]. Given the corresponding 3D coordinate $P_{\chi}\in\mathbb{R}^{N\times 3}$ from $P_{box}$, we first warp it from the 3D system to the 2D correspondence $p_{\chi}\in\mathbb{R}^{N\times 2}$ at the current time by the intrinsic and extrinsic matrix.
On the spatial hand, intending to access the target region, we sample the original feature to intermediately regress the existing offset from $p_{\chi}$ to the target. To expand the search breadth, we extend the sampling points analogical to Eq. 2 and enlarge the proportion of residual addition with a weight $\eta$. On the other temporal hand, with time going, there exists ego-car motion and object motion in the autonomous driving scenario, which requires compensation before sampling. In light of the short term in this sparse temporal stereo, i.e., $T_{\chi}$ is a small positive integer, we approximate the object motion as a uniform rectilinear motion. Thus, we first compensate $p_{\chi}$ with current velocity $P_{vel}$ and then warp it into every coordinate system as $\{p_{\chi}^{t}\}_{t=1}^{T_{\chi}}$ of per historical time by the medium transition in the global world coordinate system. Then the per-frame sampled feature $F_{\delta t},t\in\{0,1,...,T_{\chi}\}$ is formulated by
@@
$(\textbf{M}_{t}(p_{\chi}+\tau t\cdot P_{vel})+\Delta p_{\chi kt})]$
where $\tau$ is the time interval between every two adjacent frames, and $\textbf{M}_{t}$ is the ego-motion transform matrix from current time to previous $t$ time. Then the multi-frame features are fed into the sparse temporal encoder $f_{enc}$, a naive three-layer MLP, for temporal iterative fusion. Based on our approximation, the projection error will increase with $t$ larger. Thus we let $\lambda$ denote a constant in the range [0, 1], which is introduced for long-term suppression. With the iterative fusion $F_{\delta(t-1)}\leftarrow f_{enc}(\{F_{\delta(t-1)},\lambda F_{\delta t}\})$, we can get the final sparse sampled feature $F_{\delta}$ from $\{F_{\delta t}\}_{t=0}^{T_{\chi}}$. The whole box-level offset $\delta_{\chi}$ can be calculated as follows:

### v2

@@ 3 Methodology
Figure 2: Overview of LSSInst. The multi-view images with previous $T$ frames are fed into the backbone network for the image features. BEV branch looks around the image feature to generate the BEV feature by view transformation and temporal encoding. Instance adapter aggregates the sparse object-aware feature from the BEV feature and prepares the multiplicate 3D query combination. Instance branch looks back at the image feature and perfects the sparse feature by spatiotemporal sampling and fusion. Lastly, the model makes the final prediction based on the updated output.
Utilizing instance-level representations on the basis of the scene-level BEV to excavate more detailed features and geometric information is of practical significance for generalized 3D perception. In this work, we propose LSSInst, which looks back for the more geometry-aware and fine-grained target feature extraction to bridge the adaptation between scene-level and instance-level 3D representations. The overview of our framework is shown in Fig. 2, and we organize the remaining part as follows. Firstly, Sec. 3.1 briefly introduces the BEV branch. Next, Sec. 3.2 introduces the instance adapter module, and the instance branch is given in Sec. 3.3.
@@ 3.1 BEV Branch: Looking around for scene-level representation
The multi-view sequential images with the previous $T$ frames are first input into the 2D image backbone network for feature extraction. Then the BEV branch receives the extracted image feature $F_{img}\in\mathbb{R}^{N_{v}\times N_{s}\times C\times H\times W}$ and functions as an around looker, translating $F_{img}$ from 2D camera views to BEV for preliminary scene-level representation $F_{\beta}\in\mathbb{R}^{C\times H_{\beta}\times W_{\beta}}$, where $N_{v}$, $N_{s}$ denote the camera-view and scale number respectively. This branch can be briefly divided into temporally-shared view transformation for BEV generation and BEV sequence fusion. The 2D-to-BEV view transformation is naturally based on LSS paradigm, which can be mainly concluded as the depth refinement module (DRM), feature extraction net, and voxel pooling. For the best version of the framework, we adopted the Gaussian-spaced top-k stereo [34] for a better depth distribution map before voxel pooling. After the shared view transformation, a sequence of BEV representations will be aligned into current time $t$ and fed to the BEV temporal encoder to form the final current BEV. Here the encoder is designed as a very lightweight residual network for dimension reduction only.
3.2 Instance Adapter: Scene-to-instance adaptation
For the sake of preserving a coherent and solid semantic consistency between BEV and instance representations, we propose the instance adapter module to eliminate the gaps in the position description and space discrepancy. Since the BEV feature $F_{\beta}$ is a scene-level representation surrounding the ego car, there is redundancy and inflexibility in modeling instance-level features. To that end, the proposed adapter module first performs a reprojection of the proposal box coordinates $P_{o}\in\mathbb{R}^{N_{\beta}\times 3}$ obtained through the BEV proposal head, returning to the BEV-recognized position $P_{\beta}\in\mathbb{R}^{N_{\beta}\times 2}$ to resample the object-related features. Here $N_{\beta}$ denotes the number of BEV proposals. Given the BEV point-cloud range $R_{\beta}$ with the corresponding voxel size $S_{v}$ and up-sampling factor $\sigma$, we can formulate the 2D reprojected coordinate as follow:
@@ 3.2 Instance Adapter: Scene-to-instance adaptation
here $z_{\beta}$ denotes the z-axis homogeneous term and is actually a constant of 1. Moreover, due to the overfitting bias in BEV, the focused area may deviate from the actual object position. Inspired by deformable attention [56], the adapter module incorporates the learnable offsets based on the original focused feature $F_{\beta}^{{}^{\prime}}\in\mathbb{R}^{N_{\beta}\times C}$ for misalignment compensation by exploring more semantically-aware regions. Suppose $i\in\{1,2,...,N_{\beta}\}$ denotes an arbitrary element index of $F_{\beta}^{{}^{\prime}}$ and its aggregated instance-wise feature $F_{\alpha}^{i}$ can be formulated by
@@ 3.3 Instance Branch: Looking back for instance-level representation
Given the sequential image features $\{F_{img}^{t}\}_{t=0}^{T_{\chi}}$ $(T_{\chi}\leq T)$ from the image backbone network and the sparse instance features $F_{\chi}$ with corresponding 3D boxes $P_{box}\in\mathbb{R}^{N\times C_{box}}$ from the instance adapter, the instance branch will spatially and temporally look back the image feature based on the referenced box coordinates and iteratively extract the abundant but more fine-grained representations to update pre-features. This branch can be roughly regarded as a multilayer Transformer-decoder-like [42] module for 3D detection, which is briefly divided into two parts: box-level offset and embedding, as well as spatiotemporal sampling and fusion.
@@
Different from the previous DETR-like 3D approaches such as DETR3D [46], Polarformer [20], VEDet [3] which iteratively refine only with 3D coordinate offset regression, the instance branch adopts the box-level offset regression based on $P_{box}$. With this convenience, we can encode all the geometric-aware information of the entire box to substitute the transitional positional encoding, thereby expanding and enriching the space of feature expression rather than the superficial positional level. With it combined with the sparse instance features, there will be more geometric priors and implicit compensation in subsequent attention interactions. Precisely, we first categorize $P_{box}$ based on the element semantic of box dimension into four divisions, which are position $P_{pos}\in\mathbb{R}^{N\times 3}$ (i.e., $x,y,z$), scale $P_{sca}\in\mathbb{R}^{N\times 3}$ (i.e., $w,l,h$), velocity $P_{vel}\in\mathbb{R}^{N\times 2}$ (i.e., $v_{x},v_{y}$), and orientation $P_{ori}\in\mathbb{R}^{N\times 2}$ (i.e., $sin(\theta_{yaw}),cos(\theta_{yaw})$) respectively. Then we introduce five separated linear projections $\{\textbf{E}_{l3}^{j}\}^{2}_{j=1}\in\mathbb{R}^{3\times C}$, $\{\textbf{E}_{l2}^{j}\}^{2}_{j=1}\in\mathbb{R}^{2\times C}$ and $\textbf{E}_{g}\in\mathbb{R}^{C\times C}$ for comprehensive encoding, of which the former four embed every category locally and the last one embeds them globally. The final box embedding $G_{\chi}\in\mathbb{R}^{N\times C}$ can be formulated by
@@
The sparse feature $F_{\chi}$ with the box embedding $G_{\chi}$ will be updated by the spatial and temporal sampling after being fed into the multi-headed self-attention block [42]. Given the corresponding 3D coordinate $P_{\chi}\in\mathbb{R}^{N\times 3}$ from $P_{box}$, we first warp it from the 3D system to the 2D correspondence $p_{\chi}\in\mathbb{R}^{N\times 2}$ at the current time by the intrinsic and extrinsic matrix.
On the spatial hand, intending to access the target region, we sample the original feature to intermediately regress the existing offset from $p_{\chi}$ to the target. To expand the search breadth, we extend the sampling points analogous to Eq. 2 and enlarge the proportion of residual addition with a weight $\eta$. On the other temporal hand, with time going, there exists ego-car motion and object motion in the autonomous driving scenario, which requires compensation before sampling. In light of the short term in this sparse temporal stereo, i.e., $T_{\chi}$ is a small positive integer, we approximate the object motion as a uniform rectilinear motion. Thus, we first compensate $p_{\chi}$ with current 2D velocity $p_{vel}\in\mathbb{R}^{N\times 2}$, i.e., the 2D correspondence of $P_{vel}$ and then warp it into every coordinate system as $\{p_{\chi}^{t}\}_{t=1}^{T_{\chi}}$ of per historical time by the medium transition in the global world coordinate system. Then the per-frame sampled feature $F_{\delta t},t\in\{0,1,...,T_{\chi}\}$ is formulated by
@@
$(\textbf{M}_{t}(p_{\chi}-\tau t\cdot p_{vel})+\Delta p_{\chi kt})]$
where $\tau$ is the time interval between every two adjacent frames, and $\textbf{M}_{t}$ is the ego-motion transform matrix from current time to previous $t$ time. Then the multi-frame features are fed into the sparse temporal encoder $f_{enc}$, a naive three-layer MLP, for temporal iterative fusion. Based on our approximation, the projection error will increase with $t$ larger. Thus we let $\lambda$ denote a constant in the range [0, 1], which is introduced for long-term suppression. With iterative fusion $F_{\delta(t-1)}\leftarrow f_{enc}(\{F_{\delta(t-1)},\lambda F_{\delta t}\})$, we can get the final sparse sampled feature $F_{\delta}$ from $\{F_{\delta t}\}_{t=0}^{T_{\chi}}$. The whole box-level offset $\delta_{\chi}$ can be calculated as follows: